OpenStudy (anonymous):

Find the length of an arc of 40° in a circle with an 8 inch radius.

OpenStudy (campbell_st):

the circumference of a circle is \[C = 2\pi r\] the fraction of a circle you are dealing with is \[\frac{40}{360}\] so the arc length is the fraction of the circumference or \[l = \frac{40}{360} \times 2\pi r\] substitute r = 8 and calculate the answer

OpenStudy (anonymous):

What's the answer?

OpenStudy (campbell_st):

just calculate it here is a calculator if you need it http://web2.0calc.com/

OpenStudy (anonymous):

8pi/9 inches. That's what i got, is that correct?

OpenStudy (campbell_st):

I would have thought it was \[l = \frac{2\times \pi \times 8}{9}\]